40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 23:14:24 2018, aaron, Update, OMC, OMC scanning/aligning script 
    Reply  Tue Nov 20 20:33:11 2018, aaron, Update, OMC, OMC scanning/aligning script pathwaysketch.pdfIMG_0094.JPGIMG_0097.JPG
       Reply  Mon Nov 26 15:43:16 2018, aaron, Update, OMC, OMC scanning/aligning script c1omc.png
          Reply  Thu Dec 6 11:16:28 2018, aaron, Update, OMC, OMC channels 
             Reply  Mon Dec 10 12:29:05 2018, aaron, Update, OMC, OMC channels 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg545E5512-D003-408B-9F00-55F985966A16.jpegDFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
                Reply  Mon Dec 10 19:47:06 2018, aaron, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:42:44 2018, Koji, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:48:04 2018, aaron, Update, OMC, OMC channels 
                      Reply  Thu Dec 13 22:24:21 2018, aaron, Update, OMC, OMC channels 
          Reply  Mon Dec 10 12:11:28 2018, aaron, Update, OMC, Aligning the OMC B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
             Reply  Tue Dec 11 14:24:18 2018, aaron, Update, OMC, Aligning the OMC 
                Reply  Tue Dec 11 22:50:07 2018, aaron, Update, OMC, Aligning the OMC 6x
                   Reply  Thu Dec 13 22:36:42 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:02:24 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:06:12 2018, aaron, Update, OMC, Aligning the OMC 
                         Reply  Mon Dec 17 20:45:40 2018, aaron, Update, OMC, Aligning the OMC IMG_5922.JPGIMG_5921.JPG
                            Reply  Tue Dec 18 18:13:32 2018, aaron, Update, OMC, OMC L HV Piezo driver tests (again) IMG_4576_copy.jpg
                               Reply  Wed Dec 19 00:12:46 2018, gautam, Update, OMC, 40m OMC DCC node 
Message ID: 14312     Entry time: Tue Nov 20 20:33:11 2018     In reply to: 14163     Reply to this: 14317
Author: aaron 
Type: Update 
Category: OMC 
Subject: OMC scanning/aligning script 

I finished running the cabling for the OMC, which involved running 7x 50ft DB9 cables from the OMC_NORTH rack to the 1X2 rack, laying cables over others on the tray. I tried not to move other cables to the extent I could, and I didn't run the new cables under any old cables. I attach a sketch diagram of where these cables are going, not inclusive of the entire DAC/ADC signal path.

I also had to open up the AA board (D050387, D050374), because it had an IPC connector rather than the DB37 that I needed to connect. The DAC sends signals to a breakout board that is in use (D080302) and had a DB37 output free (though note this carries only 4 DAC channels). I opened up the AA board and it had two IPC 40s connected to an adapter to the final IPC 70 output. I replaced the IPC40 connectors with DB37 breakouts, and made a new slot (I couldn't find a DB37 punch, so this is not great...) on the front panel for one of them, so I can attach it to the breakout board.

I noticed there were many unused wires, so I had to confirm that I had the wiring correct (still haven't confirmed by driving the channels, but will do). There was no DCC for D080302, but I grabbed the diagrams for the whitening boards it was connected to (D020432) and for the AA board I was opening up as well as checked out elog 8814, and I think I got it. I'll confirm this manually and make a diagram if it's not fake news.

Attachment 1: pathwaysketch.pdf  26.561 MB  | Hide | Hide all
pathwaysketch.pdf
Attachment 2: IMG_0094.JPG  1.435 MB  | Hide | Hide all
IMG_0094.JPG
Attachment 3: IMG_0097.JPG  2.240 MB  | Hide | Hide all
IMG_0097.JPG
ELOG V3.1.3-